# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Appendix D Technical Details In Section 4
$\leq B\Delta T+(|\mathcal{S}|M-|\mathcal{G}|)|\mathcal{G}|(24k^{2})\frac{1}{\Delta}\log\left(\frac{e}{\Delta^{6}}\right)+\sum_{\Delta_{s}^{m}>0}(2|\mathcal{G}|+2)\Delta_{s}^{m}.$
Take $\Delta=\frac{24k^{2}|\mathcal{G}|(|\mathcal{S}|M-|\mathcal{G}|)}{BT}$ will give us the desired result. ∎
Appendix E Implementation Detail
E.1 Regret Curve
We will describe the details of implementation. Specifically, we will randomly generate a number between $[0,1]$ independently and assign this number to be $P_{a}^{m}(s,1),\forall a,m,s\in\{0,1\}$. To make sure Assumption 3.1 hold, we will switch the number if $P_{0}^{m}(s,1)>P_{1}^{m}(s,1)$.
@@ E.1 Regret Curve
Without loss of generality, assume that $I^{1}(1)\geq I^{1}(0)\geq I^{2}(1)\geq I^{2}(0)$. We choose $\gamma$ to be between $I^{1}(0)$ and $I^{2}(0)$. Specifically, let $\theta\in(0,1)$ be a tunable parameter. We set $\gamma:=\theta I^{2}(1)+(1-\theta)I^{2}(0)$. To make sure that the gap between $\gamma$ and $I^{2}(1)$ is not too small, we choose those instances where $I^{2}(1)-I^{2}(0)\geq 0.15$. To generate the three different $\gamma$ values in Figure 1, we selected three different choices of $\theta$: $0.1,0.5,0.9$, yielding the observed $\Delta_{s_{\min}}^{m_{\min}}$ values.
E.2 Cumulative Reward
@@
Here we will discuss more on the implementation of the algorithms. First of all for the UCWhittle families introduced in Wang et al. (2023b), the algorithms reinitialize the initial state after certain time-steps. We do not include this episodic structure as this contradicts our goal of maximizing the finite-horizon cumulative reward. We keep the rest structures of their algorithms as the same.
@@
We additionally note that the code provided by Wang et al. (2023b) contains errors and could not be used to reproduce the results shown in Wang et al. (2023b). A mistake was found in Line 73 of uc__\__whittle.py that they provided. Specifically, they use the current state of an agent as the index to get the current action of this agent. As a result, we build our simulation environment from scratch.
@@
Jiamin: Furthermore, we note that when the stochatsic dominance assumption does not hold, the Whittle’s index could be negative. Meanwhile, even when the stochastic dominance assumption holds, the estimator of the transition probability given by UCWhittleUCBPriv and UCWhittlePvPriv is not guaranteed to satsify the stochastic dominance assumption. Therefore, we make modifications to the binary search rountine of getting Whittle’s index. Specifically, we let the initial lower bound be $-10$ to ensure a more accurate calculation of Whittle’s index.
Adversarial Setting
@@
Figure E.8: Average cumulative reward under $B=5$ (left column), $10$ (middle column), $20$ (right column). Averaged over $50$ instance satisfying the opposite of Assumption 3.1, each with $13$ repetitions. Noiseless reward.
In this paragraph, we will modify the generated instances to make sure that not giving intervention will stochastically dominate giving intervention. After an instance is generated, we check whether there exists $s,m$ such that $P_{1}^{m}(s,1)>P_{0}^{m}(s,1)$. If there does, we will simply swap these two transitions. We will test the performance of LGGT comparing to the same sets of policies.
In Figure E.8, we notice that LGGT with threshold $\gamma=0$ performs similarly to UCWhittlePvPriv and is better than other online algorithms for all three budgets considered across 1000 horizons. We emphasize that in this setting, the optimal policy would be not giving any intervention at all, which makes the problem easy.

### v4

@@ Appendix F Technical Details In Section 4
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m^{\prime}}(s^{\prime})\leq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right].$ (F.62)
Again, for Eq. (F.61), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbbm{1}\left\{s_{\tau_{k}^{m}(s)}^{m}=s,\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\},\tau_{k}^{m}(s)\leq T\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}]\right\}\right]+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{t_{\mathrm{random}}^{k}\leq k_{3}^{k}\right\}\right],$
where $N_{s,m}^{(\tau_{k}^{m}(s))}(r):=\sum_{l=1}^{k}\mathbbm{1}\left\{a_{\tau_{l}^{m}(s)}^{m}=0,s_{\tau_{l}^{m}(s)}^{m}=s,\mathrm{random}\right\}$ denotes the total number of times that action 0 is given to agent $m$ when $m$ is in state $s$ and the algorithm is doing random exploration, $t^{k}_{\mathrm{random}}:=\sum_{l=1}^{m}\sum_{l=1}^{k}\mathbbm{1}\left\{\mathrm{random}\right\}$ denotes the total number of times when agent $m$ is in state $s$ and the algorithm is doing random explorations. We let $k_{1}=\max\left\{\frac{M-B}{2M},\frac{2M-3B}{2M}\right\},k_{2}=\min\left\{\frac{3(M-B)}{2M},\frac{2M-B}{2(M-B)},k_{3}^{k}=\frac{C\log(k)}{2\min\{\Delta^{2},1/2\}}\right\}$
For the first term, we have
$\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})t_{\mathrm{random}}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\left|\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)-I^{m}(s)\right|\geq\frac{\Delta}{2}\right\}\right]$
$\leq\sum_{k=0}^{T}\frac{1}{k^{2}}$
$\leq 2,$
where the second last inequality holds due to Lemma F.2. Similarly, we can prove the last two terms are upper bounded by 2 using Lemma F.2.
Using the same way, we can upper bound Eq. (F.62) by 6. Therefore, Eq. (F.59) can be upper bounded as
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\text{not~random}\right\}\right]$
$=\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq 12|\mathcal{G}|.$ (F.63)
Further assume that $k\left|\Delta\right|\geq\Delta_{s}^{m}$, for all $I^{m}(s)\leq\gamma$. Then combining Eq. (F.60) and Eq. (F.63) will give us the desired result.
We also have
$\mathcal{R}(\pi)$ $\leq kB\Delta T+\frac{k|\mathcal{S}|MD(M|\mathcal{S}|-|\mathcal{G}|)}{\Delta}\log T$
$\leq k\sqrt{BMD|\mathcal{S}|(M|\mathcal{S}|-|\mathcal{G}|)T\log T}.$
Appendix G Implementation Detail
All the experiments are conducted on MacBook Pro with Apple M3 Pro chip and 18 GB memory.
G.1 Regret Curve
Figure G.8: Regret Curve as horizon $T$ grows.
To numerically verify Theorem 4.4, we set $M=2$, $B=1$, and $\mathcal{S}=\{0,1\}$. We set the expected reward of the state as $R^{m}(s)=s$. We randomly generated $10$ instances and chose three $\gamma$ values satisfying Assumption 4.2, labeling each $\gamma$ by the corresponding gap $\Delta$ between the threshold and the smallest incremental reward. Specifically, we will randomly generate a number between $[0,1]$ independently and assign this number to be $P_{a}^{m}(s,1),\forall a,m,s\in\{0,1\}$.
@@ G.1 Regret Curve
Without loss of generality, assume that $I^{1}(1)\geq I^{1}(0)\geq I^{2}(1)\geq I^{2}(0)$. We choose $\gamma$ to be between $I^{1}(0)$ and $I^{2}(0)$. Specifically, let $\theta\in(0,1)$ be a tunable parameter. We set $\gamma:=\theta I^{2}(1)+(1-\theta)I^{2}(0)$. To make sure that the gap between $\gamma$ and $I^{2}(1)$ is not too small, we choose those instances where $I^{2}(1)-I^{2}(0)\geq 0.15$. To generate the three different $\gamma$ values in Figure G.8, we selected three different choices of $\theta$: $0.1,0.5,0.9$, yielding the observed $\Delta$ values.
For each $\gamma$, the regret was averaged over $100$ repetitions per instance (Figure G.8). Figure G.8 demonstrates sublinear regret for all choices of $\gamma$, with regret stabilizing at a constant as the horizon length increases, consistent with Theorem 4.4. Further, as $\Delta$ decreases, regret increases as it becomes harder to distinguish whether the incremental reward is above the threshold or not. This matches the results in Theorem 4.4.
G.2 Cumulative Reward
We first describe the benchmarks we chose in detail:
  • Oracle Whittle: A variant of Whittle’s index policy with access to the ground truth transitions and rewards, optimized for maximizing discounted reward (as described in Appendix B).
  • OMR: A modified LP-based policy for maximizing finite-horizon reward.
  • UCWhittle-penalty (Wang et al., 2023b): It constructs the confidence interval of the unknown transition probability and then chooses one to calculate Whittle’s index policy. Specifically, it constructs a confidence interval and chooses the transition probability that will maximize $V_{\lambda}^{m}(s)$ (Eq. (B.9)) for some adaptively updated $\lambda$.
  • RRMABB: It uses the explore-then-commit framework. It first randomly gives action and constructs the confidence interval of the unknown transition. After the exploration period, it uses the estimated transition to calculate the OMR policy.
In all of the settings, the discount factor $\beta$ was set to $0.9$. We note that though $\beta=1$ was chosen in Biswas et al. (2021), $\beta=0.9$ has better empirical performance, which aligns with the choice in Wang et al. (2023b).
@@ G.2 Cumulative Reward
Then, we will discuss more on the implementation of the algorithms. First of all for the UCWhittle families introduced in Wang et al. (2023b), the algorithms reinitialize the initial state after certain time-steps. We do not include this episodic structure as this contradicts our goal of maximizing the finite-horizon cumulative reward. We keep the rest structures of their algorithms as the same.
@@
We additionally note that the code provided by Wang et al. (2023b) contains errors and could not be used to reproduce the results shown in Wang et al. (2023b). A mistake was found in Line 73 of uc_\_whittle.py that they provided. Specifically, they use the current state of an agent as the index to get the current action of this agent. As a result, we build our simulation environment from scratch.
@@
Furthermore, we note that the Whittle’s index could be negative. Therefore, we make modifications to the binary search routine of getting Whittle’s index. Specifically, we let the initial lower bound be 10-10 to ensure a more accurate calculation of Whittle’s index.
We also note that in the original RRMABB algorithm, it assumes that they can reinitialize at any state for arbitrary number of times. This makes their exploration criteria, which will only stop when sufficient number of samples are collected for each agent, state, action pair, possible. We think this assumption is not impractical and do not use this structure. Therefore, we also change the algorithm so that the exploration period will end after the same number of time-steps instead of number of samples.
Choice of $\gamma$
In this paragraph, we will compare $\epsilon$-GT with different threshold with 2 states. Specifically, we will choose $\gamma\in\{-0.9,0.9\}$. 0.9-0.9 denotes a more moderate choice of $\gamma$ which balances between satisfying Assumption 4.2 and maximizing the reward and $0.9$ denotes a more aggressive choice of $\gamma$ which aims at maximizing the reward. As shown in Figure G.9, $\gamma=-0.9$ outperforms $\gamma=0.9$ across all instances.
@@
Figure G.9: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Two states.
Randomly Generated Instances With 10 States
Figure G.10: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Ten states.
In this paragraph, we increase the size of the state space to 10 states and do not impose the stochastic dominance assumption to randomly generate 50 instances to test the performance of $\epsilon$-GT comparing to the benchmarks. Similarly, we also test two settings: 1) a noiseless case where the reward is deterministic, and 2) a noisy case where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 0.1. We would like to note that since UCWhittlePv and RRMABB behave like random policy for 2 states, we do not include that in the 10 states setting.
As shown in Figure G.10, $\epsilon$-GT with threshold $\gamma=-8.9$ consistently outperform WIQL in both noiseless and noisy setting. Further, we would like to emphasize that the gap between $\epsilon$-GT and WIQL is larger than the 2 states setting, demonstrating that our algorithm works well in larger states setting and adapting to the real-life setting better.
G.3 Ablation Study
In this section, we will test the performance of Alg. 1 with different choices of $\eta$. Specifically, we will choose four choices of $\eta\in\{0.1,0.5,1,1.5\}$. We will use the same experiment setup as in Section 5 and fix $\gamma=-0.9$.
Figure G.11: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Two states.
As shown in Figure G.11, Alg. 1 is stable over different choices of $\eta$.